OpenStudy (anonymous):

While showing off for some grade nine girls at the ski hill, Eero wipes out on a snowboard. After the crash he tumbled 30m in 4.2s, what was his speed in: a) m/s b) km/hr?

OpenStudy (anonymous):

hi

OpenStudy (cggurumanjunath):

30/4.2 m/s

OpenStudy (cggurumanjunath):

speed=distance in m/time in sec

OpenStudy (anonymous):

what would the answer be

OpenStudy (cggurumanjunath):

how much is 30/4.2 ? @thecat229

OpenStudy (anonymous):

7.14

OpenStudy (cggurumanjunath):

yes

OpenStudy (cggurumanjunath):

7.14 m/s

OpenStudy (anonymous):

ok thanks

OpenStudy (cggurumanjunath):

1m/s=3.6km/hr 7.14m/s=7.14*3.6

OpenStudy (cggurumanjunath):

how much is 7.14*3.6 ? @thecat229